Calculate the volume in liters of a 0.0743/molL potassium dichromate solution that contains 150.g of potassium dichromate K2Cr2O7

Round your answer to 3 significant digits.
Answer #1

Moles of K2CrO7 = mass / molar mass = 150 / 294.18 = 0.510 moles

Also, Moles of K2CrO7 = molarity * Volume

0.510 = 0.0743 * V

Solving, V = 6.864 Liters